OpenStudy (anonymous):

Please I need help this... what is the solution of 6x + y = 24 and 3x + y/3 =18 ? a. x=10, y=-36 b. x=24, y= -1 c. x=-12, y=-15 d. x=18, y=1 e. x=-1, y=60

OpenStudy (haseeb96):

y=24-6x put it in the second equation and get the y and after getting y value then put y value in any equation and get x

OpenStudy (haseeb96):

A is the correct answer but u should solve u it

OpenStudy (anonymous):

6(10) + -36 = 24. Use a calculator and see if that works. 3(10) + -36/3 =18 ? What about this one? Yes. They both work. The answer is A. Just try plugging in numbers, if you get the right answer then you've solved the problem.

OpenStudy (anonymous):

ah okay thanks for the help you guys

OpenStudy (haseeb96):

3x+(24-6x) /3 =18 and now taking common 3 3x + 3 (8 -2x) /3 =18 so the 3 and 3 cancel to eachother 3x +8 -2x =18 x=18-8 x=10 now put x into the first equation 6(10) + y =24 60 + y = 24 y=24 -60 y=-36 is the correct answer i think u have got it how to solve this type of question

OpenStudy (haseeb96):

@melanielovee4 good job Girl !!! :)